nathan is p years now and in 10 years time he will be 3 times as old as martin. express martins age 10 years from now in terms of p
Found 2 solutions by algebrahouse.com, Boreal:
Answer by algebrahouse.com(1659)

p = Nathan's age now
p + 10 = Nathan's age in 10 years {added 10 onto current age}

In 10 years, Nathan will be 3 times as old as Martin.

p + 10
------- = Martin's age in 10 years {Nathan's age divided by 3}
3

To ask me a question, visit: www.algebrahouse.com



Answer by Boreal(15235)

Nathan is p
Martin is m
right now they are p and m years old
in 10 years, they will be P+10 and M+10
P+10=3(M+10)=3M+30
3M=P-20
M=(1/3)(P-20)
NOTE: I am assuming Martin is m years old now. If that is not the case, then Nathan is P, and in 10 years he will be P+10, and that is 3M
then M=(1/3)(P+10)
The question needs to be clarified. Many of these age problems start from a given year and both are looked at in some time in the future.